# MemRL: Self-Evolving Agents via Runtime Reinforcement Learning on Episodic Memory

### v1

@@ 4.5 Stability Analysis
We analyze the stability of MemRL from a reinforcement learning perspective, focusing on the convergence behavior of the utility estimates stored in memory. Unlike classical value iteration, MemRL performs non-parametric runtime learning using a constant-step-size update. We show that under mild and realistic assumptions, the learned utility values converge in expectation to stable estimates of memory effectiveness, with bounded variance.
@@
At each time step $t$, the agent observes an intent state $s_{t}$, retrieves a memory item $m_{t}\in\mathcal{M}_{t}$, generates an output $y_{t}$, and receives a scalar reward $r_{t}\in[-1,1]$ indicating task success or failure. The generation policy follows the decomposition defined in Eq. 1, where $\mu$ denotes the retrieval policy and $p_{\mathrm{LLM}}$ is a frozen inference policy.
@@
Moreover, the convergence rate is exponential:
@@
The stability of the local estimate (Theorem 1) extends to the global memory utility $Q(m)$. By the linearity of expectation, $Q(m)$ acts as a Monte Carlo integrator striving to converge to:
@@
A theoretical challenge arises here: the weighting term $\Pr(s|m)$ is a latent variable governed by the retrieval policy $\mu(m|s;\mathcal{M})$, which itself shifts as $Q$-values evolve. To prove convergence despite this dependency, we analyze MemRL as a Generalized Expectation-Maximization (GEM) process (dempster1977maximum; neal1998view). From a variational perspective, the system performs coordinate ascent on a global objective function $\mathcal{J}(Q,\mu)$ (the variational lower bound of expected reward): (i) E-Step (Policy Improvement): The Phase-B ranking updates the retrieval policy $\mu$ to align with current estimates, monotonically increasing $\mathcal{J}$ with respect to $\mu$; (ii) M-Step (Value Update): The utility update (Eq. 8) increases $\mathcal{J}$ with respect to $Q$. By the Monotonic Improvement Theorem (neal1998view), this alternating optimization guarantees that the system converges to a stationary point where the retrieve policy stabilizes ($\mu_{t+1}\approx\mu_{t}$). Consequently, the induced distribution $\Pr(s|m)$ becomes time-invariant, ensuring that Eq. 13 holds and effectively preventing catastrophic forgetting by anchoring updates to a stable policy. More details can be found in Appendix B.
5 Experiments
5.1 Experimental Setup
Baselines.
We evaluate MemRL against a comprehensive suite of memory-augmented baselines under a unified frozen-backbone setting to isolate the contribution of memory mechanisms: (i) RAG-based Approaches: RAG (lewis2020retrieval) and Self-RAG (asai2023selfrag), representing standard semantic retrieval and critique-based filtering, respectively. (ii) Agentic Memory: Mem0 (chhikara2025mem0) and MemP (fang2025memp), which introduce structured write/read operations or procedural memory distillation. (iii) Test-Time Scaling: Pass@k and Reflexion, which applies iterative self-refinement over $k$ rounds (shinn2023reflexion).
Benchmarks.
We evaluate MemRL and baselines on four diverse benchmarks: BigCodeBench (zhuo2025bigcode) for code generation, ALFWorld (shridhar2021alfworld) for embodied navigation, Lifelong Agent Bench (zheng2025lifelong) for OS/DB interaction, and Humanity’s Last Exam (HLE) (phan2025hle) for multidisciplinary complex reasoning.
Table 1: Runtime Learning main results. We compare MemRL against various baselines over 10 epochs. The results are reported as Last Epoch Accuracy / Cumulative Success Rate (CSR). CSR indicates the percentage of tasks solved at least once during the training process. “–” indicates the experiment was not applicable.
BigCodeBench Lifelong Agent Bench ALFWorld HLE
Method
Code Gen
(Last / CSR)
OS Task
(Last / CSR)
DB Task
(Last / CSR)
Exploration
(Last / CSR)
Knowledge Frontier
(Last / CSR)
Model GPT-4o GPT-4o-mini GPT-4o-mini GPT-4o-mini Gemini-3-pro
No Memory 0.485 0.674 0.860 0.278 0.357
Pass@10 – / 0.577 – / 0.756 – / 0.928 – / 0.462 – / 0.524
Reflexion – / 0.614 – / 0.714 – / 0.938 – / 0.358 – / –
RAG 0.475 / 0.483 0.690 / 0.700 0.914 / 0.916 0.370 / 0.415 0.430 / 0.475
Self-RAG 0.497 / 0.561 0.646 / 0.732 0.891 / 0.898 0.290 /0.290 0.411 / 0.427
Mem0 0.487 / 0.495 0.670 / 0.702 0.920 / 0.926 – / – 0.400 / 0.406
MemP 0.578 / 0.602 0.736 / 0.742 0.960 / 0.966 0.324 / 0.456 0.528 / 0.582
MemRL (ours) 0.595 / 0.627 0.794 / 0.816 0.960 / 0.972 0.507 / 0.697 0.573 / 0.613
  • $\dagger$
    Experiments are still running, results are reported using fewer than 10 training epochs.
We evaluate our MemRL and baselines under two distinct settings: Runtime Learning, which assesses the ability to learn and adapt within a training session, and Transferring, which evaluates the generalization capability of the learned memory on unseen tasks.
Runtime Learning Results.
Table 1 reports the Last Epoch Accuracy (Last) and Cumulative Success Rate (CSR) over 10 training epochs. MemRL consistently outperforms all baselines across all domains, validating that non-parametric value estimation effectively guides self-evolution. Notably, the advantages of our value-aware retrieval are most pronounced in exploration-heavy environments such as ALFWorld. Here, MemRL achieves a remarkable last-epoch accuracy of 0.507, representing a relative improvement of approximately 56% over MemP (0.324) and 82% over the No Memory baseline (0.278). Furthermore, the high CSR of 0.697 in ALFWorld indicates that the RL component effectively encourages the agent to explore and discover solutions for complex tasks that similarity-based retrieval methods often fail to solve. Similarly, in the challenging Knowledge Frontier HLE benchmark, MemRL improves the last accuracy to 0.573 compared to 0.528 for MemP, where the CSR even reached an astonishing 61.3%. Comparing these gains against single-turn tasks like BigCodeBench reveals an important trend: the performance uplift correlates with task structural complexity. The benefits of MemRL are maximized in environments characterized by deep exploration and high procedural transferability (e.g., ALFWorld), whereas the margin is narrower in tasks with lower structural reuse. This suggests that our value-based mechanism is particularly adept at distilling and transferring complex problem-solving patterns from exploratory trajectories.
Overall, the simultaneous improvement in both CSR and Last Accuracy suggests that MemRL not only discovers high-quality solutions during exploration but also effectively retains and retrieves them for stable performance.
Table 2: Transfer Learning results on BigCodeBench, Lifelong Agent Bench and ALFWorld. We compare our MemRL against various retrieval and memory baselines using best validation results.
BigCodeBench Lifelong Agent Bench ALFWorld
Method
Code Generation
OS Task
DB Task
Exploration
Model GPT-4o GPT-4o-mini GPT-4o-mini GPT-4o-mini
No Memory 0.485 0.673 0.841 0.314
RAG 0.479 0.713 0.920 0.336
Self-RAG 0.500 0.653 0.881 0.293
Mem0 0.485 0.686 0.935
MemP 0.494 0.720 0.928 0.421
MemRL (ours) 0.508 0.746 0.942 0.479
  • $\dagger$
    Experiments are still running, results are reported using fewer than 10 training epochs.
Transferring Results.
We evaluate memory transferability by freezing the memory bank after training and testing on held-out sets (30% split). As shown in Table 2, MemRL exhibits superior generalization compared to baseline methods. On BigCodeBench, MemRL achieves the highest accuracy of 0.508, outperforming advanced retrieval methods such as Self-RAG (0.500) and standard MemP (0.494). In the OS control tasks (Lifelong Agent Bench), our method attains an accuracy of 0.746, significantly improving upon the standard RAG baseline (0.713). Consistent with the runtime results, the gain is substantial in ALFWorld, where MemRL reaches 0.479, demonstrating a clear margin over MemP (0.421) and RAG (0.336). These results validate that the Value-Aware Retrieval and Non-Parametric RL mechanism in MemRL does not merely overfit to training instances; instead, it filters out low-value memories, retaining high-utility experiences that facilitate generalization to unseen tasks.
5.2 Ablations
5.2.1 Effectiveness of Runtime RL
(a) Success Rate by Epoch
(b) Cumulative Success Rate (CSR)
Figure 5: OS Interaction Performance. (a) MemRL demonstrates superior stability and higher peak performance compared to the baseline. (b) The widening gap in CSR illustrates that RL-enhanced methods effectively accumulate solved tasks over time.
To isolate the effect of runtime RL, we compare the memory mechanisms with and without the RL component (i.e., MemP vs. MemRL and RAG vs. MemRL(RAG)) within the OS interaction environment. Figure 5(a) illustrates the success rate evolution across epochs. While the RL enhancement yields marginal gains in the initial phase, a clear performance divergence emerges as training progresses. MemRL consistently outperforms the vanilla MemP baseline in later epochs, exhibiting a smoother learning curve with fewer regressions. This stability suggests that the RL-driven value function effectively mitigates the interference of noisy memories that often plague static similarity-based retrieval.
We further analyze the CSR in Figure 5(b), which measures the agent’s ability to solve distinct tasks at least once during the training process. Here, the benefits of RL are even more pronounced. MemRL and MemRL(RAG) attain final CSR outperforming their counterparts MemP and RAG, respectively. The monotonic widening of the performance gap in Figure 5(a) indicates that runtime RL significantly enhances the agent’s ability to recover from early failures. By prioritizing memories with high expected utility, the agent effectively mitigates the interference of noisy memories and consolidates successful experiences, transforming transient exploration into robust, repeatable capabilities.
5.2.2 Impact of Q-Value Weighting
To understand the interplay between semantic retrieval and reinforcement learning, we analyze the impact of the Q-weighting factor $\lambda$ in the scoring function, i.e., the Equation 6: We compare the balanced setting ($\lambda=0.5$) against two extremes: pure semantic retrieval ($\lambda=0$) and pure greedy RL ($\lambda=1$).
(a) Success Rate
(b) Cumulative Success Rate
Figure 6: Ablation study on Q-value weighting factor $\lambda$. The balanced setting ($\lambda=0.5$) achieves superior stability and peak performance compared to extreme configurations.
As illustrated in Figure 6, the balanced configuration consistently yields superior performance. This indicates that the combination of semantic grounding and utility-driven ranking effectively guides the agent towards high-quality solutions while maintaining context relevance. Besides, the pure semantic retrieval baseline provides a stable starting point but lacks the mechanism to filter out suboptimal memories that are semantically similar but functionally incorrect. Consequently, its performance plateaus early, which highlights the necessity of the RL signal for continued improvement. On the other hand, the pure RL setting reveals the risks of disregarding semantic context, exhibiting significant instability and poor initial performance. We attribute this instability to context detachment: without the constraint of semantic similarity, the agent may retrieve high-$Q$ memories (successful in other contexts) that are irrelevant to the current task. This confirms that semantic similarity must act as an anchor for retrieval, while RL provides the necessary gradient for optimizing memory selection.
5.2.3 Sensitivity to Retrieval Size ($k_{1}$ and $k_{2}$).
Figure 7: Ablation on Retrieval Size ($k_{1},k_{2}$). Performance (acc: Epoch-Acc; cum: Cumulative Acc) comparison on HLE (CS/AI) across different retrieval counts.
To investigate the impact of memory capacity on reasoning performance, we conducted an ablation study on a subset of the HLE benchmark (Computer Science/AI category). We compared two retrieval configurations: a larger recall setting ($k_{1}=10,k_{2}=5$) and a compact recall setting ($k_{1}=5,k_{2}=3$). As shown in Figure 7, the compact setting ($k_{1}=5,k_{2}=3$) achieves superior stability compared to the larger setting, suggesting that for complex reasoning tasks like HLE, simply increasing context volume can introduce irrelevant noise that interferes with the model’s judgment. Therefore, a smaller, higher-quality set of retrieved memories is sufficient for maintaining reasoning precision.
5.3 Discussion
In this section, we delve deeper into the mechanisms driving MemRL’s performance, connecting empirical results back to the Stability-Plasticity framework proposed in Section 4.
5.3.1 MemRL as a Trajectory Verifier.
Benchmark Interaction MemP (%) MemRL (%) Gain (pp)
ALFWorld Multi-step 45.6 69.7 +24.1
OS Task Multi-step 74.3 81.7 +7.4
HLE Single-step 58.2 61.3 +3.1
BigCodeBench Single-step 60.2 62.7 +2.5
Table 3 reveals a correlation between task structural complexity and performance gain. The gains are most profound in multi-step sequential tasks (e.g., ALFWorld +24.1 Percentage Points(pp)) compared to single-turn tasks (e.g., BigCodeBench +2.5 pp). In sequential tasks, a retrieved memory must be valid for the entire trajectory. Standard semantic retrieval often fetches memories that match the initial instruction but fail in later steps. By propagating the final reward backward to the memory utility $Q$, MemRL effectively learns to verify the whole trajectory, filtering out brittle policies that look correct only on the surface.
5.3.2 Predictive Power of the Q Critic
(a) Success Rate vs. Q-Range
(b) Memory Composition
Does the learned Q-value truly reflect solution quality? Figure 8(a) shows a strong positive correlation (Pearson $r=0.861$) between the Critic’s estimated Q-values and empirical task success rates. The success rate increases from 21.5% in the lowest-confidence bin to 88.1% in the highest, indicating that a major driver of performance is the Critic’s ability to rank memories by their likelihood of leading to successful task completion.
Further analysis of memory composition (Figure 8(b)) suggests a secondary source of gain: robustness. Even in high-Q bins (0.90.91.01.0), the agent retains a small fraction of memories labeled as “failure” ($\sim$12%). Rather than contradicting the Critic, this pattern is consistent with the interpretation that Q-values can capture utility beyond binary outcomes: some unsuccessful trajectories remain strategically useful because they encode near-correct reasoning and transferable procedural lessons.
Case study: a high-Q “failure” memory as a transferable near-miss.
We find high-Q failure memories that summarize a minor, localized mistake and a corrective heuristic that generalizes across tasks. Here we use near-miss to denote trajectories that are largely correct but fail due to minor, localized errors (e.g., verification slips or tool-usage details). For example, one failure memory with $Q=0.9878$ corresponds to a trajectory that followed a correct approach but incorrectly treated an empty command output as evidence of failure. The stored reflection explicitly identifies the root cause (misinterpreting empty output), warns against the pattern (equating “no output” with failure), and recommends the correct approach (validate success via exit status, error logs, or other objective signals). When retrieved in later episodes, this single “failure” memory supports perfect downstream outcomes in our logs (15/15 successes), demonstrating that the memory is valuable precisely because it captures a fixable near-miss rather than an uninformative failure.
Taken together, these results suggest that the Critic is not merely separating “success” from “failure,” but assigning higher value to memories that provide reusable guidance—including a subset of high-utility failures that are near-miss in nature. By retaining and reusing such corrective heuristics, the agent can be more robust than simple success-replay mechanisms.
5.3.3 Stability of MemRL
We analyze the underlying mechanisms of MemRL through the view of the stability-plasticity dilemma, examining how our framework balances the acquisition of new capabilities with the retention of established knowledge.
The superior CSR (Table 1) demonstrates that MemRL effectively expands the agent’s solution space. Unlike heuristic baselines constrained by static similarity, which often retrieve redundant nearest neighbors leading to repeated failures, MemRL allows the agent to identify and reinforce non-obvious but effective strategies. This capability enables the agent to break through local optima where traditional retrieval methods stagnate.
Figure 9: Epoch Acc and Cumulative ACC of MemRL and MemP in HLE.
Besides, long-term training dynamics (Figure 9) reveal a critical stability advantage. Heuristic methods like MemP suffer from a widening gap between CSR and current epoch accuracy, indicating that new explorations inadvertently overwrite effective historical policies (catastrophic forgetting). In contrast, MemRL maintains synchronized growth. We attribute this to our theoretical guarantee. From a general MDP perspective (Eq. 3), the value update benefits from the standard Bellman contraction, $\|\mathcal{T}Q-Q^{*}\|_{\infty}\leq\gamma\|Q-Q^{*}\|_{\infty}$ (sutton2018reinforcement), shrinking the error by $\gamma$ at each step. More specifically, under our Monte Carlo style modeling (Eq. 8), this process is guaranteed by Section 4.5. Unlike heuristic ranking which may drift randomly, our approach mathematically constrains the policy to climb the variational lower bound of expected reward, ensuring the stable, non-decreasing performance observed in our experiments.
We further validate these insights using the Forgetting Rate, defined as the ratio of tasks that regress from success in the previous round to failure in the current round ($\texttt{success}\to\texttt{fail}$). As shown in Figure 11, MemRL achieves the lowest mean forgetting rate ($0.041$), outperforming the baseline MemP ($0.051$) and empirically confirming our analysis.
The Necessity of Normalization and Similarity Gate. Incidentally, the Figure 11 also highlights the necessity of our stabilization design: removing normalization and lowering similarity (w/o Norm_SimGate) threshold causes the mean forgetting rate to spike to $0.073$ due to unconstrained utility variance. This demonstrates that z-score normalization and strict similarity gating are essential for filtering noise, ensuring that the self-evolution remains stable while maximizing positive transfer.
Figure 10: Forgetting Rate in HLE
Figure 11: Similarity-based Generalization
5.3.4 Impact of Task Similarity on Memory Efficacy.
To understand the underlying conditions where MemRL thrives, we analyze the correlation between the intra-dataset semantic similarity ($\text{Sim}_{intra}$) and the absolute performance gain provided by our method ($\Delta=\text{Acc}_{\text{MemRL}}-\text{Acc}_{\text{NoMem}}$).
As illustrated in Figure 11, we analyze the correlation between intra-dataset semantic similarity and the absolute performance gain ($\Delta$) provided by MemRL. The linear regression trend reveals a general positive correlation: environments with higher structural repetition allow the agent to retrieve and reuse optimal policies more effectively. At the upper extreme, ALFWorld (similarity $0.518$) acts as a strong anchor point for this trend, exhibiting the highest repetition and a corresponding maximum performance boost ($\Delta=+0.229$). This confirms that for highly repetitive procedural tasks, memory serves as an effective shortcut to optimal trajectories. Following the regression line, benchmarks with moderate similarity—such as Lifelong-OS ($0.390$) and BigCodeBench (0.3080.308)—cluster in the middle region, showing steady improvements ($\Delta\approx+0.11\sim+0.12$) where the agent successfully generalizes coding patterns or OS commands across related instructions.
The HLE Anomaly: Generalization vs. Memorization.
HLE presents a unique outlier. Despite having the lowest similarity ($0.186$) due to its diverse, multi-disciplinary nature, it exhibits a surprisingly high runtime gain ($0.357\rightarrow 0.573,\Delta=+0.216$). This gain operates on a different mechanism than ALFWorld. In high-similarity benchmarks, MemRL succeeds via Positive Transfer—generalizing shared patterns to new instances. In contrast, the gain in HLE stems from Runtime Memorization. Since HLE questions are distinct and domain-specific, the agent relies on the Runtime Learning phase to ”memorize” specific solutions to difficult problems through repeated exposure. This distinction highlights MemRL’s versatility: it supports both pattern generalization in structured domains and specific knowledge acquisition in diverse domains.
6 Conclusion
In this paper, we introduced MemRL, a novel framework that enables LLMs to self-evolve through non-parametric reinforcement learning on episodic memory. Addressing the limitations of semantic retrieval and the instability of parameter fine-tuning, MemRL treats memory retrieval as a value-based decision process. By structuring memory into Intent-Experience-Utility triplets and applying Bellman updates, the agent learns to differentiate high-value strategies from semantic noise without modifying the backbone model weights.
Our extensive evaluations across diverse domains—ranging from code generation to embodied navigation—demonstrate that MemRL significantly outperforms existing memory-augmented baselines in both runtime learning and generalization to unseen tasks. Theoretical and empirical analyses further reveal that MemRL effectively resolves the stability-plasticity dilemma: the frozen LLM provides robust reasoning, while the evolving memory utility acts as a plastic channel for adaptation. Moreover, we identify that the utility-driven retrieval mechanism functions as a trajectory verifier, enabling the agent to filter out brittle policies in complex, multi-step tasks. We hope this work establishes a new paradigm for building self-evolving agents that can continuously learn from interaction in a stable and efficient manner.
Appendix A Theoretical Analysis and Proofs
In this section, we provide the detailed derivation for the convergence of the Q-value estimation under the Exponential Moving Average (EMA) update rule, and extend the analysis to the global stability of memory utility under task distributions.
A.1 Proof of Theorem 1: Convergence of EMA Estimation
@@
    Stationary Reward. The reward $r_{t}$ at step $t$ is drawn from a distribution with a constant mean $\beta(s,m)=\mathbb{E}[r_{t}|s,m]$ and finite variance $\sigma^{2}$.
@@
We define $\mathcal{F}_{t}$ as the filtration (history) up to time $t$. Taking the conditional expectation of Eq. 14 given $\mathcal{F}_{t}$:
@@
A.2 Bounded Variance and Global Stability
@@
Appendix B Theoretical Analysis: Convergence via Variational Inference
@@ B.1 The Convergence Objective
B.2 Variational Objective with Trust Region
@@
Here, the first term directly corresponds to the expected utility $\mathbb{E}[Q_{t}(m)]$ we aim to converge, while $\pi_{\text{sim}}$ represents the fixed semantic prior (derived from Phase-A). The KL-divergence term acts as a regularizer crucial for two reasons:
@@
B.3 Optimization via Generalized Expectation-Maximization (GEM)

### v2

@@ A.1 Stability Analysis
We analyze the stability of MemRL from a reinforcement learning perspective, focusing on the convergence behavior of the utility estimates stored in memory. MemRL performs non-parametric runtime learning using a constant-step-size update. We show that under mild and realistic assumptions, the learned utility values converge in expectation to stable estimates of memory effectiveness, with bounded variance.
@@
At each time step $t$, the agent observes an intent state $s_{t}$, retrieves a memory item $m_{t}\in\mathcal{M}_{t}$, generates an output $a_{t}$, and receives a scalar reward $r_{t}\in[-1,1]$ indicating task success or failure. The generation policy follows the decomposition defined in Eq. 1, where $\mu$ denotes the retrieval policy and $p_{\mathrm{LLM}}$ is a frozen inference policy.
@@
where the expectation is taken over the stochastic rewards resulting from the Inference Policy $p_{LLM}$ distribution.
@@
Moreover, the convergence rate is exponential (Sutton and Barto, 2018):
@@
The stability of the local estimate (Theorem A.1) extends to the global memory utility $Q(m)$. By the linearity of expectation, $Q(m)$ acts as a Monte Carlo integrator striving to converge to:
@@
A theoretical challenge arises here: the weighting term $\Pr(s|m)$ is a latent variable governed by the retrieval policy $\mu(m|s;\mathcal{M})$, which itself shifts as $Q$-values evolve. To prove convergence despite this dependency, we analyze MemRL as a Generalized Expectation-Maximization (GEM) process (Dempster et al., 1977; Neal and Hinton, 1998). From a variational perspective, the system performs coordinate ascent on a global objective function $\mathcal{J}(Q,\mu)$ (the variational lower bound of expected reward): (i) E-Step (Policy Improvement): The Phase-B ranking updates the retrieval policy $\mu$ to align with current estimates, monotonically increasing $\mathcal{J}$ with respect to $\mu$; (ii) M-Step (Value Update): The utility update (Eq. 4) increases $\mathcal{J}$ with respect to $Q$. By the Monotonic Improvement Theorem (Neal and Hinton, 1998), this alternating optimization guarantees that the system converges to a stationary point where the retrieve policy stabilizes ($\mu_{t+1}\approx\mu_{t}$). Consequently, the induced distribution $\Pr(s|m)$ becomes time-invariant, ensuring that Eq. 14 holds and effectively preventing catastrophic forgetting by anchoring updates to a stable policy. More details can be found in Appendix A.4.
A.2 Proof of Theorem A.1: Convergence of EMA Estimation
@@
    Stationary Reward. The reward $r_{t}$ at step $t$ is drawn from a distribution induced by the stochastic action generation $a\sim p_{LLM}(a_{t}|s_{t},m)$, with a constant mean $\beta(s,m)=\mathbb{E}[r_{t}|s,m]$ and finite variance $\sigma^{2}$.
@@
We define $\mathcal{F}_{t}$ as the filtration (history) up to time $t$. Since the reward $r_{t}$ depends on the action $a_{t}$ sampled subsequently from the frozen LLM, taking the conditional expectation of Eq. 15 given $\mathcal{F}_{t}$:
@@
A.3 Bounded Variance and Global Stability
@@
A.4 Convergence via Variational Inference
@@ A.4.1 The Convergence Objective
A.4.2 Variational Objective with Trust Region
@@
Here, the expectation is taken over the state distribution $\mathcal{D}$. The first term directly corresponds to the expected utility $\mathbb{E}[Q_{t}(m)]$ we aim to converge, while $\pi_{\text{sim}}$ represents the fixed semantic prior (derived from Phase-A). The KL-divergence term acts as a regularizer crucial for two reasons:
@@
A.4.3 Optimization via Generalized Expectation-Maximization (GEM)